I am trying to help my daughter Ember in her algebra class.
How would you solve and simplify the following expression: 16 1/4

Answer by stanbon(75887)

How would you solve and simplify the following expression: 16 1/4
----
16 1/4 is a mixed number.
---
YOu can convert it to an improper fration as follows:
-------
16 1/4 = (4*16+1)/4 = (64+1)/4 = 65/4
============================
Cheers,
Stan H.